# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 6. A gluing technique for the $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ case
$G=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})$
From (6.8), we can construct a section map
$\psi:G\rightarrow B^{2880},$
such that $\pi_{q_{1},q_{2}}(\psi(x))=x$.
Since $q_{3}^{\prime}<q_{3}^{40\theta}$, there is $q_{4}||q_{3},q_{4}>q_{3}^{1-\theta^{\frac{1}{2}}}$ such that for every $p^{n}||q_{4}$, we have $p^{[40\theta^{\frac{1}{2}}n]}\nmid q_{3}^{\prime}$.
Write $q_{4}=\prod_{j\in J}p_{j}^{n_{j}}$. For each $p_{j}^{n_{j}}||q_{4}$, we consider $\psi_{j}=\pi_{p_{j}^{[n_{j}\theta^{1/4}]}}\circ\mathbb{P}_{1}\circ\psi$.
Define
According to Proposition 3.1, there are two scenarios:
  1. (1)
    $|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (6.10)
  2. (2)
    There is a subset $S_{j}\in G$, $|S_{j}|\geq\frac{99}{100}|G|$ such that $\psi_{j}=h_{j}$ over $S_{j}$ where $h_{j}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]})$.
Let $J=J_{1}\sqcup J_{2}$ where $J_{1},J_{2}$ is the collection of indices falling into Case (1) and Case (2), respectively. Write $q_{4}=q_{4}^{\prime}q_{4}^{\prime\prime}$, where
$q_{4}^{\prime}=\prod_{j\in J_{1}}p_{j}^{n_{j}},\hskip 28.45274ptq_{4}^{\prime\prime}=\prod_{j\in J_{2}}p_{j}^{n_{j}}.$
We further write $q_{4}^{\prime\prime}=q_{5}q_{5}^{\prime}$, where
$q_{5}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}=1\end{subarray}}p_{j}^{n_{j}},\hskip 28.45274ptq_{5}^{\prime}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}\neq 1\end{subarray}}p_{j}^{n_{j}}.$
We divide our analysis into three cases.
6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
In this case we have
$\sum_{j\in J_{1}}(\log p_{j}^{n_{j}})|\mathcal{G}_{j}|>\log\left((q_{4}^{\prime})^{10^{-4}}\right)|G|^{2}$ (6.11)
The left side of (6.11) is equal to
$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(\prod_{j\in U}p_{j}^{n_{j}})|\cap_{j\in U}\mathcal{G}_{j}\bigcap\cap_{j\in J_{1}-U}\mathcal{G}_{j}^{c}|$ (6.12)
Since the number of subsets of $J_{1}$ is $<q_{4}^{0+}$, there exists $J_{1}^{\prime}\subset J_{1}$, $\tilde{q}=\prod_{i\in J_{1}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime})^{\frac{1}{2}10^{-4}}$, such that
$|\cap_{i\in J_{1}}\mathcal{G}_{i}|>q^{0-}|G|^{2}.$
Take any $(g_{1},g_{2})\in\cap_{i\in J_{1}^{\prime}}\mathcal{G}_{i}$, and consider $\gamma_{0}=\psi(g_{1})\psi(g_{2})\psi(g_{1}g_{2})^{-1}$. Then $\gamma_{0}$ satisfies,
$\pi_{q_{1},q_{2}}(\gamma_{0})=1,$
and for any $p^{n}||\tilde{q}$,
$\pi_{p^{[n\theta^{{1}/{4}}]}}\circ\mathbb{P}_{1}(\gamma_{0})\neq 1.$
Take a set $B^{\prime}\subset B^{2880}$ which maps bijectively to $\Gamma(q_{3}^{\prime})/\Gamma(q_{3})$ under the projection $\pi_{q_{3},1}$. Take
$B^{\prime\prime}=\{b_{1}\gamma_{0}b_{1}^{-1}\gamma_{0}^{-1}\cdot(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})\gamma_{0}(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})^{-1}\gamma_{0}^{-1},b_{1},b_{2}\in B^{\prime}\}$
It follows from a quantitative open mapping result due to Salehi-Golsefidy (see Propostions 49, Proposition 49’ of [Gol19]) that
$\pi_{q_{3}}\circ\mathbb{P}_{1}\left(B^{\prime\prime}\right)\supset\Gamma(q_{3}^{\{O(\theta^{1/4})\}}).$ (6.13)
by first taking the logarithmic map to work on the Lie algebra level, then taking the exponential map back. The definition for $B^{\prime\prime}$ is motivated by the fact that for any $v\in\mathfrak{sl}_{2}(\mathbb{Q})$, $\mathfrak{sl}_{2}(\mathbb{Q})=[\mathfrak{sl}_{2}(\mathbb{Q}),v]+[[\mathfrak{sl}_{2}(\mathbb{Q}),v],v]$, where $+$ here means set addition.
Obviously, $\pi_{q_{1},q_{2}}(B^{\prime\prime})=1$. Multiplying $B^{\prime\prime}$ with $B^{2880}$ and observing (6.8) we obtain Proposition 6.1 in this case by setting $q_{3}^{*}=\tilde{q}$.
6.2. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
The local homomorphisms $h_{j},j\in J_{2}$ can be lifted to a homomorphism
$h:G\rightarrow\Gamma(\prod_{j\in J_{2}}p_{j})/\Gamma((q_{4}^{\prime\prime})^{\{\theta^{\frac{1}{4}}\}}).$
Following the previous reasoning for obtaining $J_{1}^{\prime}$, there is a set $J_{2}^{\prime}\subset J_{2}$, such that
$\bar{q}=\prod_{i\in J_{2}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime\prime})^{\frac{99}{200}},$
$|\cap_{i\in J_{2}}S_{i}|>(q_{1}q_{2})^{-\theta}|G|,$
and $\psi\equiv h$ on $S=\cap_{i\in J_{2}}S_{i}$.
By Proposition 3.12, we have $S^{2880}\supset G^{\prime}=\Gamma(q_{1}^{\prime\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime\prime})/\Gamma(q_{2})$ for some $q_{1}^{\prime\prime}||q_{1}$, $q_{2}^{\prime\prime}||q_{2}$, $q_{1}^{\prime\prime}q_{2}^{\prime\prime}<(q_{1}q_{2})^{80\theta}$. This implies the existence of a subgroup $G^{\prime\prime}$ of $G^{\prime}$ of the form
$G^{\prime\prime}=\Gamma(\tilde{q}_{1}(q_{1}^{*})^{\{\theta^{\frac{1}{2}}\}})/\Gamma(q_{1})\times\Gamma(\tilde{q}_{2}(q_{2}^{*})^{\{\theta^{{1}/{2}}\}})/\Gamma(q_{2}),$ (6.14)
where
$q_{1}=\tilde{q_{1}}q_{1}^{*},q_{2}=\tilde{q_{2}}q_{2}^{*},\tilde{q}_{1},q_{1}^{*}||q_{1},\tilde{q}_{2},q_{2}^{*}||q_{2},\tilde{q}_{1}\tilde{q}_{2}<(q_{1}q_{2})^{80\theta^{{1}/{2}}}.$ (6.15)
Then we can define a map $\tilde{\psi}:G^{\prime\prime}\rightarrow B^{2880^{2}}$ as, for any $x\in G^{\prime\prime}$, choose a word $x=s_{1}s_{2}\cdots s_{2880}$ and let
$\tilde{\psi}(x)=\psi(s_{1})\psi(s_{2})\cdots\psi(s_{2880}).$ (6.16)
Then we have $\pi_{(q_{4}^{\prime\prime})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=h(x)$, and thus $\pi_{(q_{5})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=1$.
Recall Lemma 5.43. By iteratively taking commutator of $\tilde{\psi}(G^{\prime\prime})$ $O(\log(1/\theta))$ times, we recover a group
$\Gamma((q_{1}^{*})^{\{{2\theta^{{1}/{4}}}\}}q_{5})/\Gamma(q_{1}^{*}q_{5})\times\Gamma((q_{2}^{*})^{\{{2\theta^{{1}/{4}}}\}})/\Gamma(q_{2}^{*})\subset\pi_{q_{1}^{*}q_{5},q_{2}^{*}}(B^{O(\log(1/\theta))}),$ (6.17)
which is identity when projected to $q_{5}$ and is large when projected to $(q_{1},q_{2}^{\prime})$.
Recall $q_{5}||q_{3}$ and $v_{p}(q_{3}^{\prime})<v_{p}(q_{3})$ for every $p^{n}||q_{5}$. It follows from (6.15), (6.17), and (6.9) that
$|\pi_{q_{1}q_{5},q_{2}}(B^{O(\log 1/\theta)})|>(q_{1}q_{2}q_{5})^{3-{O(\theta^{{1}/{4}})}}.$ (6.18)
In this case, we take $q_{3}^{*}=q_{5}$ and the proposition is proved.
6.3. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}^{\prime}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
We begin with a lemma.
Lemma 6.19.
Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$. Then $p_{1}=p_{2}$ and $p^{x}||\xi-1$ for some $x\leq m_{1}+n_{2}-m_{2}$.
Proof.
If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$ and we obtaine a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$. Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor thorough the center $Z$. $f^{\prime}$ can not be injective because there are elements in $\Gamma_{p_{1}}$ of order $p_{1}$, which is coprime to $p_{2}$. Likewise, $Z$ can not be the kernel of $f^{\prime}$ because one can also find an element $x\in\Gamma_{p_{1}}/Z$ such that the order of $x$ is coprime to $p_{2}$. Hence, $f^{\prime}$ as well as $f$ must be trivial and we arrive at a contradiction, so $p_{1}=p_{2}$.
Suppose $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$. Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $x\leq m_{1}+n_{2}-m_{2}$. ∎
Lemma 6.19 implies that $q_{5}^{\prime}||q_{2}^{*}$, since $(q_{5}^{\prime},q_{1}^{*})=1$.
Let $G^{\prime\prime}$ and $\widehat{\psi}$ be the same as in Case 6.2.
The following isomorphism is clear
$G^{\prime\prime}\cong\Big{(}\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\Big{)}\times\Big{(}\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})\Big{)}\times\Big{(}\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\Big{)}$
Since $\pi_{(q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}}\circ\mathbb{P}_{1}\circ\tilde{\psi}$ is a homomorphism and factors through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, it induces a homomorphism
$\widehat{\psi}:\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\rightarrow\Gamma/\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}).$
Since $\widehat{\psi}$ is nontrivial after reduction by each exact divisor of $(q_{5}^{\prime})^{\left\{\frac{1}{2}\theta^{1/4}\right\}}$, we can find $g_{1}\in\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}})/\Gamma(q_{5}^{\prime})$ such that for each $p|q_{5}^{\prime}$,
$v_{p}(g_{1})=O(n\theta^{1/4}),\hskip 14.22636ptv_{p}(\widehat{\psi}(g_{1}))<(\frac{1}{2}n\theta^{1/4}).$
Let $g_{2}=g_{1}^{p^{\{\frac{1}{4}n\theta^{1/4}\}}}$, then
$v_{p}(g_{2})=\Theta(n\theta^{1/4}),\hskip 14.22636pt\frac{1}{4}n\theta^{1/4}<v_{p}(\widehat{\psi}(g_{2}))<\frac{3}{4}n\theta^{1/4}.$
Write $g_{2}\equiv 1+Q_{1}X_{1}(\operatorname{mod}Q_{1}^{2}),\widehat{\psi}(g_{2})\equiv 1+Q_{2}X_{2}(\operatorname{mod}Q_{2}^{2})$ for some primitive $X_{1},X_{2}\in\text{Lie}(\text{SL}_{2}(\mathbb{Z}))$. Collecting all powers of $g_{2}$, pulling them back to $G^{\prime\prime}$ and using $\widehat{\psi}$ is induced from $\pi_{(q_{5}^{\prime})^{\theta^{1/4}}}\circ\mathbb{P}_{1}\circ\psi$ by factoring through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, we obtain
$\pi_{q_{1}^{*}Q_{2}^{\{{4}/{3}\}},(q_{2}^{*}/q_{5}^{\prime})Q_{1}^{2}}(B^{2880^{2}})\cong\left(\pi_{q_{1}^{*},q_{2}^{*}/q_{5}^{\prime}}\times\pi_{Q_{2}^{\{4/3\}},Q_{1}^{2}}\right)(B^{2880^{2}})$
$\supset$ $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma(({q_{2}^{*}}/{q_{5}^{\prime}})^{\{\theta^{1/2}\}})/\Gamma({q_{2}^{*}}/{q_{5}^{\prime}})\times\left(\pi_{Q_{2}^{\{{4}/{3}\}},Q_{1}^{2}}\left((1,1)+\mathbb{Z}(Q_{2}X_{2},Q_{1}X_{1})\right)\right)$ (6.20)
with $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta^{1/4})$.
With (6.3) at hand, which plays the same role as Theorem B creating a one-parameter group, we run the same arguments in Section 5.5, and we give a sketch here. We first apply Proposition 4.2 with $Q=q_{5}^{\prime}$ to find elements from $A$ to conjugate $X=(X_{1},X_{2})$ to other directions. By considering five primitive linear forms, we can produce $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ such that the following holds: There exists $\bar{q_{5}}||q_{5}^{\prime}$ and
$\bar{q_{5}}>(q_{5}^{\prime})^{1-5\theta^{\frac{1}{2}}},$ (6.21)
such that
$(\bar{q_{5}})^{\{5\theta^{\frac{1}{2}}\}}\pi_{\bar{q_{5}}}(V\times V)\subset\pi_{\bar{q_{5}}}\text{Span}_{\mathbb{Z}}\{X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}\}$ (6.22)
as long as $\delta<c_{2}\eta\theta^{1/2}/2$.
Multiplying the RHS of (6.3) with its conjugates by $g_{1},g_{2},g_{3},g_{4},g_{5}$, followed by taking commutators of this product iteratively, and finally multiplying all these commutators together, we obtain
$\Gamma((q_{1}^{*}\bar{q_{5}}))^{\{5\theta^{1/2}\}}/\Gamma(q_{1}^{*}\bar{q_{5}})\times\Gamma(((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})^{\{5\theta^{1/2}\}})/\Gamma((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})$
$\subset$ $\pi_{q_{1}^{*}\bar{q_{5}},(q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}}}\{B\cup A\}^{O(\log 1/\theta)}$
and thus
$|\pi_{q_{1}{q_{5}^{\prime}},q_{2}}\{B\cup A\}^{O(\log 1/\theta)}|>(q_{1}^{*}q_{2}^{*}\bar{q_{5}})^{3(1-5\theta^{1/2})-}>(q_{1}q_{2}q_{5}^{\prime})^{3-O(\theta^{1/2})}$
In this case, we let $q_{3}^{*}={q_{5}^{\prime}}$ and the proposition is proved.
@@ 6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
Let $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ to be determined at (7.12). Let $A$ be given in Proposition 2.2 but we assume
$|\pi_{q}(A\cdot A\cdot A)|<|A|^{1+\delta}.$ (7.1)

### v4

@@ 6.1. The case $q^{J_{1}}\geq(q_{3}^{\prime})^{\frac{1}{2}}$
$\bar{\psi}(g)\equiv(1+Q_{5}X_{2},1+Q_{4}X_{1},)(\operatorname{mod}Q_{5}^{2},Q_{4}^{2}),$
with $X_{1},X_{2}\in\text{Mat}_{2}{(\mathbb{Z})}$ primitive and traceless, $Q_{4}|q^{J_{6}}$, $Q_{5}|q^{J_{6}}$. From (6.33) and (6.34), we have
$v_{p_{j}}(Q_{4})=\Theta(n_{i}\theta^{\frac{1}{4}})$
$\frac{1}{4}n_{i}\theta^{\frac{1}{4}}\leq v_{p_{i}}(Q_{5})\leq\frac{3}{4}n_{i}\theta^{\frac{1}{4}},\hskip 14.22636pt\forall i\in J_{6}.$
Let $Q_{6}=\text{gcd}(Q_{5}^{2},{({q^{J_{6}}})^{\{\theta^{\frac{1}{4}}\}}})$, recalling that $\mathbb{P}_{3}\circ\bar{\psi}$ is a homomorphism up to reduction by $({q^{J_{6}}})^{\{\theta^{\frac{1}{4}}\}}$}.
We consider the family
$\Xi=\{\bar{\psi}(g^{n}):n\in\mathbb{Z}\}\subset B_{4}.$ (6.35)
Each $\bar{\psi}(g^{n})$ satisfies
$\mathbb{P}_{1}(\bar{\psi}(g^{n}))=1$
$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(\bar{\psi}(g^{n}))=1,\hskip 14.22636pt\forall i\in I_{2}-J_{6}$
$\bar{\psi}(g^{n})\equiv(1+nQ_{5}X_{2},1+nQ_{4}X_{1})(\operatorname{mod}Q_{6},Q_{4}^{2})$
with
$v_{p_{j}}(Q_{4}),v_{p_{j}}(Q_{5})=\Theta(n_{j}\theta^{\frac{1}{4}})$
and
$\frac{1}{3}v_{p_{j}}(Q_{5})\leq v_{p_{j}}(Q_{6})-v_{p_{j}}(Q_{5})\leq v_{p_{j}}(Q_{5}),\hskip 14.22636pt\forall j\in J_{6}.$
With $\Xi$ at hand, which plays the same role as the one-parameter group given by Lemma 5.50, we can run the same arguments in §5.5 and §5.6. Recall $\delta<\theta$. We start by applying Proposition 4.2 to find $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ to conjugate $X=(X_{2},X_{1})$ so that
$\text{gcd}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}),q^{J_{6}})=(q^{J_{6}})^{O(\theta)},$
which implies a set $J_{7}\subset J_{6}$, such that
$q^{J_{7}}>(q^{J_{6}})^{1-O(\theta^{\frac{1}{2}})}$
$v_{p_{j}}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}))<n_{j}\theta^{\frac{1}{2}},\forall j\in J_{7}.$
Then we apply the argument §5.6 to create a set $B_{5}\subset\{B_{4}\cup A\}^{O(\log\frac{1}{\theta})}$ such that
$|\pi_{q^{J_{7}},q^{J_{7}}}(B_{5})|>(q^{J_{7}})^{6-O(\theta^{\frac{1}{4}})},$
recalling $\pi_{q^{J_{7}},q^{J_{7}}}$ is the reduction map from $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ to $\Gamma_{q^{J_{7}}}\times\Gamma_{q^{J_{7}}}$. Therefore, $B_{5}$ has the following property:
$\mathbb{P}_{1}(B_{5})=1$
$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(B_{5})=1,\forall i\in I_{2}-J_{6}$
$|\pi_{q^{J_{6}},q^{J_{6}}}(B_{5})|\geq|\pi_{q^{J_{7}},q^{J_{7}}}(B_{5})|>(q^{J_{7}})^{6-O(\theta^{\frac{1}{4}})}=(q^{J_{6}})^{6-O(\theta^{\frac{1}{4}})}.$
So there must be some element $x_{0}\in\Gamma_{q^{J_{6}}}$ such that $B_{x_{0}}:=\{g\in B_{5}:\pi_{q^{J_{6}}}\circ\mathbb{P}_{2}(g)=x_{0}\}$ satisfies
$|\mathbb{P}_{1,2}(B_{x_{0}})|=1$
$|\pi_{q^{J_{6}}}\circ\mathbb{P}_{3}(B_{x_{0}})|=(q^{J_{6}})^{3-O(\theta^{\frac{1}{4}})}.$
Then $|\pi_{q_{1}q^{J_{6}},q_{2}}(B_{1}\cdot B_{x_{0}})|=|\mathbb{P}_{1,2}(B_{1})||\pi_{q^{J_{6}}}\circ\mathbb{P}_{3}(B_{x_{0}})|>(q_{1}q_{2}q^{J_{6}})^{3-O(\theta^{\frac{1}{4}})}$. Proposition 6.1 is thus proved in this case as well with $q_{3}^{*}=q^{J_{6}}$. ∎
@@
We free all the notations newly introduced in §5 and §6 except those in the statements of Proposition 5.1 and Proposition 6.1. We first prove Proposition 2.2 for all $q=\prod_{i\in I}p_{i}^{n_{i}}$ with $n_{i}>L,\forall i\in I$, where $L>0$ depends only on $S$ and $\varepsilon$ and is determined at (7.4). Suppose $A\subset\Lambda$ satisfies (2.3), but fails (2.4), that is, $A$ satisfies
$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta}.$ (7.1)